# Online Variational Sequential Monte Carlo

### v1

@@ A.4. Auxiliary results
(71) $\left\lvert\beta_{\theta}^{N}\prod_{k=1}^{N}a_{\theta}^{k}-\beta_{\theta^{\prime}}^{N}\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\lVert\theta-\theta^{\prime}\rVert\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N}).$
@@
(72) $\left\lvert T_{\theta}h(z_{t})-T_{\theta^{\prime}}h(z_{t})\right\rvert\leq\|h\|_{\infty}\idotsint\bar{m}(x_{t},x_{t+1})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})\\ \times(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\lVert\theta-\theta^{\prime}\rVert\int_{x_{t}^{1:N}}\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N})(d\tilde{x}_{t}^{1:N})\,\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N})\\ =(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$
@@
(73) $\left\lvert T_{\theta}^{t+1}h(z)-T_{\theta^{\prime}}^{t+1}h(z)\right\rvert$ $\leq\sum_{s=0}^{t}\int T_{\theta^{\prime}}^{t-s}(z,dz^{\prime})|T_{\theta}\tilde{T}_{\theta}^{s}h(z^{\prime})-T_{\theta^{\prime}}\tilde{T}_{\theta}^{s}h(z^{\prime})|~$
(74) $\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert\sum_{s=0}^{t}\varrho^{s}$
(75) $\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}(1-\varrho)}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$
@@
(76) $|\tilde{T}_{\theta}^{t+1}h(z)-\tilde{T}_{\theta^{\prime}}^{t+1}h(z)|$ $=\left\lvert\sum_{s=0}^{t}\iint\tilde{T}_{\theta}^{s}h(z^{\prime\prime})(T_{\theta}-T_{\theta^{\prime}})(z^{\prime},dz^{\prime\prime})\tilde{T}_{\theta^{\prime}}^{t-s}(z,dz^{\prime})\right\rvert$
(77) $\leq\sum_{s=0}^{t}\left|~\int\tilde{T}_{\theta^{\prime}}^{t-s}(z,dz^{\prime})(T_{\theta}\tilde{T}_{\theta}^{s}h(z^{\prime})-T_{\theta^{\prime}}\tilde{T}_{\theta}^{s}h(z^{\prime}))\right|$
(78) $\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\lVert\theta-\theta^{\prime}\rVert\sum_{s=0}^{t}\varrho^{t-s}\|\tilde{T}_{\theta^{\prime}}^{s}h(\cdot)\|_{\infty}$
(79) $\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}(t+1)\varrho^{t}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$
@@
(80) $\left\lvert\tau_{\theta}h-\tau_{\theta^{\prime}}h\right\rvert$ $\leq|T_{\theta}^{t}h(z)-T_{\theta^{\prime}}^{t}h(z)|+|\tilde{T}_{\theta}^{t}h(z)|+|\tilde{T}_{\theta^{\prime}}^{t}h(z)|$
(81) $\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}(1-\varrho)}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert+2\varrho^{t}\|h\|_{\infty},$
@@
(82) $\left\lvert\tau_{\theta}h-\tau_{\theta^{\prime}}h\right\rvert\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}(1-\varrho)}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$
@@
(83) $\varsigma\coloneqq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\max\left\{\sup_{t\in\mathbb{N}}t\varrho^{t/2-1},\frac{1}{1-\varrho}\right\}.$
@@
(84) $\lVert H_{\theta}(z)\rVert\leq\frac{\sum_{i=1}^{N}\lVert\nabla w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)\rVert}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}^{\ell},f_{\theta}(\tilde{x}^{\ell},y,v^{\ell}),y)}\leq\frac{\tilde{\kappa}}{\epsilon^{3}},$
from which (i) immediately follows. We turn to (ii). Using Lemma A.12(i) and (84), for every $t\in\mathbb{N}$ and $z\in\mathsf{Z}$,
(85) $\lVert\mathbb{E}_{\theta}[H_{\theta}(Z_{t}^{\theta})\mid Z_{0}^{\theta}=z]-h(\theta)\rVert=\lVert T_{\theta}^{t}H_{\theta}(z)-h(\theta)\rVert=\lVert\tilde{T}_{\theta}^{t}H_{\theta}(z)\rVert\leq\frac{\tilde{\kappa}}{\epsilon^{3}}\varrho^{t}.$
@@
which proves (ii). In order to establish (iii), let
(86) $\tilde{H}_{\theta}(z)\coloneqq\sum_{s=0}^{\infty}(T_{\theta}^{s}H_{\theta}(z)-h(\theta)),\quad z\in\mathsf{Z}.$
@@
(87) $\|\tilde{H}_{\theta}(z)\|\leq\sum_{s=0}^{\infty}\lVert T_{\theta}^{s}H_{\theta}(z)-h(\theta)\rVert\leq\frac{\tilde{\kappa}}{\epsilon^{3}}\sum_{s=0}^{\infty}\varrho^{s}=\frac{\tilde{\kappa}}{\epsilon^{3}(1-\varrho)}.$
@@
(88) $T_{\theta}\tilde{H}_{\theta}(z)=\sum_{s=1}^{\infty}(T_{\theta}^{s}H_{\theta}(z)-h(\theta)),$
implying (iii). To prove (iv), write, using Lemma A.11,
(89) $\lVert H_{\theta}(z)-H_{\theta^{\prime}}(z)\rVert$ $\leq\frac{\sum_{i=1}^{N}\lVert\nabla w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)-\nabla w_{\theta^{\prime}}(\tilde{x}^{i},f_{\theta^{\prime}}(\tilde{x}^{i},y,v^{i}),y)\rVert}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}^{\ell},f_{\theta}(\tilde{x}^{\ell},y,v^{\ell}),y)}$
(90) $\hskip 56.9055pt+\frac{\lVert H_{\theta^{\prime}}(z)\rVert\sum_{i=1}^{N}\left\lvert w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)-w_{\theta^{\prime}}(\tilde{x}^{i},f_{\theta^{\prime}}(\tilde{x}^{i},y,v^{i}),y)\right\rvert}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}^{\ell},f_{\theta}(\tilde{x}^{\ell},y,v^{\ell}),y)}$
(91) $\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\lVert\theta-\theta^{\prime}\rVert,$
@@
(92) $\lVert\tilde{T}_{\theta}^{t}H_{\theta}(z)-\tilde{T}_{\theta^{\prime}}^{t}H_{\theta^{\prime}}(z)\rVert$ $\leq\lVert\tilde{T}_{\theta}^{t}H_{\theta}(z^{\prime})-\tilde{T}_{\theta}^{t}H_{\theta^{\prime}}(z^{\prime})\rVert-\lVert\tilde{T}_{\theta}^{t}H_{\theta^{\prime}}(z)-\tilde{T}_{\theta^{\prime}}^{t}H_{\theta^{\prime}}(z)\rVert$
(93) $\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\varrho^{t}\lVert\theta-\theta^{\prime}\rVert+\frac{\tilde{\kappa}}{\epsilon^{3}}\varsigma\varrho^{t/2}\lVert\theta-\theta^{\prime}\rVert$
(94) $\leq 2\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\varsigma\varrho^{t/2}\lVert\theta-\theta^{\prime}\rVert.$
@@
(95) $\lVert T_{\theta}\tilde{H}_{\theta}(z)-T_{\theta^{\prime}}\tilde{H}_{\theta^{\prime}}(z)\rVert$ $=\left\|\sum_{s=1}^{\infty}(T_{\theta}^{s}H_{\theta}(z)-h(\theta))-\sum_{s=1}^{\infty}(T_{\theta^{\prime}}^{s}H_{\theta^{\prime}}(z)-h(\theta^{\prime}))\right\|$
(96) $\leq\sum_{s=0}^{\infty}\lVert\tilde{T}_{\theta}^{s}H_{\theta}(z)-\tilde{T}_{\theta^{\prime}}^{s}H_{\theta^{\prime}}(z)\rVert$
(97) $\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\frac{2\varsigma}{1-\sqrt{\varrho}}\lVert\theta-\theta^{\prime}\rVert.$
@@
(98) $\lVert h(\theta)-h(\theta^{\prime})\rVert\leq\int\lVert H_{\theta}(z)-H_{\theta^{\prime}}(z)\rVert\,\tau_{\theta}(dz)+\|\tau_{\theta}H_{\theta^{\prime}}-\tau_{\theta^{\prime}}H_{\theta^{\prime}}\|~\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)(1+\varsigma)\lVert\theta-\theta^{\prime}\rVert,$
@@
Appendix B Details on the deep generative model of Section 6.3
In this section we provide more details and experimental results on the model presented in Section 5.3.
B.1. Data generation
@@ B.2. Model architecture
(99) $p(x_{1:T},h_{0:T},y_{1:T})=p_{0}(h_{0})\prod_{t=0}^{T-1}m_{\theta}(x_{t+1}\mid h_{t})g_{\theta}(y_{t+1}\mid h_{t},x_{t+1})p_{\lambda,\theta}(h_{t+1}\mid h_{t},x_{t+1},y_{t+1}),$
@@
(100) $H_{0}$ $\sim N(0,I),$
(101) $X_{t+1}\mid H_{t}=h_{t}$ $\sim N(\mu_{\theta}^{x}(h_{t}),\sigma_{\theta}^{x}(h_{t})^{2}),$
(102) $Y_{t+1}\mid H_{t}=h_{t},X_{t+1}=x_{t+1}$ $\sim\operatorname{Bernoulli}(\mu_{\theta}^{y}(\varphi_{\theta}^{x}(x_{t+1}),h_{t})),$
(103) $H_{t+1}\mid H_{t}=h_{t},X_{t+1}=x_{t+1},Y_{t+1}=y_{t+1}$ $\sim\delta_{\texttt{GRU}_{\lambda}(h_{t},\varphi_{\theta}^{x}(x_{t+1}),\varphi_{\lambda}^{y}(y_{t+1}))},$
@@
(104) $X_{t+1}\mid Y_{t+1}=y_{t+1},H_{t}=h_{t}\sim N(\mu_{\lambda}^{p}(\varphi_{\lambda}^{y}(y_{t+1}),h_{t}),\sigma_{\lambda}^{p}(\varphi_{\lambda}^{y}(y_{t+1}),h_{t})^{2}).$
@@
19:return $(\xi_{t-38:t+1}^{x,i},\xi_{t-39:t}^{h,i},\omega_{t+1}^{i})_{i=1}^{N},\theta_{t+1},\lambda_{t+1}$
B.3. Sampling from the generative model
In Section 5.3 of the main paper we have displayed the learning curve of the incremental ELBO through a $10^{5}$ frames-long video. The goal of this section is to analyze the quality of the learned generative model by visualizing some independently sampled videos and comparing them with some reference one generated from a different seed than the one used for training. We consider a 150 frames-long reference video, and for each sample we use the proposal distribution to generate the latent states for the first 60 iterations, i.e., we access the true frame to sample the states and then decode them, then after $t=60$ we proceed using the generative model only. Figures 710 display every third frame for the true video and five independent sample videos. We observe that the first 60 frames look similar to the true one for all samples, as expected since the model has access to the true observations; after that, the sample videos continue similarly in an initial phase, but start to move with different paces when the agent disappears behind the rectangle after $t=75$ (Figure 9). Except for some inconsistencies—such as the agent coming out from above instead of below the rectangle in the second sample of Figure 9—suggesting that the model might be further improved, we may confirm that the model predicts smooth sequential behaviors of the agent.
Figure 7. Comparison between original video and five samples under the learned model. Until $t=60$, frames are generated using the learned proposal and the given data to encode the latent states; after this, the frames are produced using the generative model. Only every $3$rd frame is displayed.
Figure 8. Comparison between original video and five samples under the learned model. Until $t=60$, frames are generated using the learned proposal and the given data to encode the latent states; after this, the frames are produced using the generative model. Only every $3$rd frame is displayed.
Figure 9. Comparison between original video and five samples under the learned model. Until $t=60$, frames are generated using the learned proposal and the given data to encode the latent states; after this, the frames are produced using the generative model. Only every $3$rd frame is displayed.
Figure 10. Comparison between original video and five samples under the learned model. Until $t=60$, frames are generated using the learned proposal and the given data to encode the latent states; after this, the frames are produced using the generative model. Only every $3$rd frame is displayed.

### v3

@@ A.4. Auxiliary results
$\left\lvert\beta_{\theta}^{N}\prod_{k=1}^{N}a_{\theta}^{k}-\beta_{\theta^{\prime}}^{N}\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\lVert\theta-\theta^{\prime}\rVert\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N}).$ (69)
@@
$\left\lvert T_{\theta}h(z_{t})-T_{\theta^{\prime}}h(z_{t})\right\rvert\leq\|h\|_{\infty}\idotsint\bar{m}(x_{t},x_{t+1})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})\\ \times(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\lVert\theta-\theta^{\prime}\rVert\int_{x_{t}^{1:N}}\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N})(d\tilde{x}_{t}^{1:N})\,\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N})\\ =(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$ (70)
@@
$\left\lvert T_{\theta}^{t+1}h(z)-T_{\theta^{\prime}}^{t+1}h(z)\right\rvert$ $\leq\sum_{s=0}^{t}\int T_{\theta^{\prime}}^{t-s}(z,dz^{\prime})|T_{\theta}\tilde{T}_{\theta}^{s}h(z^{\prime})-T_{\theta^{\prime}}\tilde{T}_{\theta}^{s}h(z^{\prime})$ | UNKNOWN (71)
$\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert\sum_{s=0}^{t}\varrho^{s}$ (72)
$\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}(1-\varrho)}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$ (73)
@@
$|\tilde{T}_{\theta}^{t+1}h(z)-\tilde{T}_{\theta^{\prime}}^{t+1}h(z)|$ $=\left\lvert\sum_{s=0}^{t}\iint\tilde{T}_{\theta}^{s}h(z^{\prime\prime})(T_{\theta}-T_{\theta^{\prime}})(z^{\prime},dz^{\prime\prime})\tilde{T}_{\theta^{\prime}}^{t-s}(z,dz^{\prime})\right\rvert$ (74)
[…]
$\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\lVert\theta-\theta^{\prime}\rVert\sum_{s=0}^{t}\varrho^{t-s}\|\tilde{T}_{\theta^{\prime}}^{s}h(\cdot)\|_{\infty}$ (76)
$\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}(t+1)\varrho^{t}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$ (77)
@@
$\left\lvert\tau_{\theta}h-\tau_{\theta^{\prime}}h\right\rvert$ $\leq|T_{\theta}^{t}h(z)-T_{\theta^{\prime}}^{t}h(z)|+|\tilde{T}_{\theta}^{t}h(z)|+|\tilde{T}_{\theta^{\prime}}^{t}h(z)|$ (78)
$\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}(1-\varrho)}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert+2\varrho^{t}\|h\|_{\infty},$ (79)
@@
$\left\lvert\tau_{\theta}h-\tau_{\theta^{\prime}}h\right\rvert\leq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}(1-\varrho)}\|h\|_{\infty}\lVert\theta-\theta^{\prime}\rVert.$ (80)
@@
$\varsigma\coloneqq(\epsilon^{6}\tilde{\kappa}+\tilde{\kappa}+\epsilon^{4}\kappa)\frac{N}{\epsilon^{7N+3}}\max\left\{\sup_{t\in\mathbb{N}}t\varrho^{t/2-1},\frac{1}{1-\varrho}\right\}.$ (81)
@@
$\lVert H_{\theta}(z)\rVert\leq\frac{\sum_{i=1}^{N}\lVert\nabla w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)\rVert}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}^{\ell},f_{\theta}(\tilde{x}^{\ell},y,v^{\ell}),y)}\leq\frac{\tilde{\kappa}}{\epsilon^{3}},$ (82)
from which (i) immediately follows.
We turn to (ii). Using Lemma A.12(i) and (82), for every $t\in\mathbb{N}$ and $z\in\mathsf{Z}$,
$\lVert\mathbb{E}_{\theta}[H_{\theta}(Z_{t}^{\theta})\mid Z_{0}^{\theta}=z]-h(\theta)\rVert=\lVert T_{\theta}^{t}H_{\theta}(z)-h(\theta)\rVert=\lVert\tilde{T}_{\theta}^{t}H_{\theta}(z)\rVert\leq\frac{\tilde{\kappa}}{\epsilon^{3}}\varrho^{t}.$ (83)
@@
which proves (ii).
In order to establish (iii), let
$\tilde{H}_{\theta}(z)\coloneqq\sum_{s=0}^{\infty}(T_{\theta}^{s}H_{\theta}(z)-h(\theta)),\quad z\in\mathsf{Z}.$ (84)
@@
$\|\tilde{H}_{\theta}(z)\|\leq\sum_{s=0}^{\infty}\lVert T_{\theta}^{s}H_{\theta}(z)-h(\theta)\rVert\leq\frac{\tilde{\kappa}}{\epsilon^{3}}\sum_{s=0}^{\infty}\varrho^{s}=\frac{\tilde{\kappa}}{\epsilon^{3}(1-\varrho)}.$ (85)
@@
$T_{\theta}\tilde{H}_{\theta}(z)=\sum_{s=1}^{\infty}(T_{\theta}^{s}H_{\theta}(z)-h(\theta)),$ (86)
implying (iii).
To prove (iv), write, using Lemma A.11,
$\lVert H_{\theta}(z)-H_{\theta^{\prime}}(z)\rVert$ $\leq\frac{\sum_{i=1}^{N}\lVert\nabla w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)-\nabla w_{\theta^{\prime}}(\tilde{x}^{i},f_{\theta^{\prime}}(\tilde{x}^{i},y,v^{i}),y)\rVert}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}^{\ell},f_{\theta}(\tilde{x}^{\ell},y,v^{\ell}),y)}$ (87)
$\hskip 56.9055pt+\frac{\lVert H_{\theta^{\prime}}(z)\rVert\sum_{i=1}^{N}\left\lvert w_{\theta}(\tilde{x}^{i},f_{\theta}(\tilde{x}^{i},y,v^{i}),y)-w_{\theta^{\prime}}(\tilde{x}^{i},f_{\theta^{\prime}}(\tilde{x}^{i},y,v^{i}),y)\right\rvert}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}^{\ell},f_{\theta}(\tilde{x}^{\ell},y,v^{\ell}),y)}$ (88)
$\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\lVert\theta-\theta^{\prime}\rVert,$ (89)
@@
$\lVert\tilde{T}_{\theta}^{t}H_{\theta}(z)-\tilde{T}_{\theta^{\prime}}^{t}H_{\theta^{\prime}}(z)\rVert$ $\leq\lVert\tilde{T}_{\theta}^{t}H_{\theta}(z^{\prime})-\tilde{T}_{\theta}^{t}H_{\theta^{\prime}}(z^{\prime})\rVert-\lVert\tilde{T}_{\theta}^{t}H_{\theta^{\prime}}(z)-\tilde{T}_{\theta^{\prime}}^{t}H_{\theta^{\prime}}(z)\rVert$ (90)
$\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\varrho^{t}\lVert\theta-\theta^{\prime}\rVert+\frac{\tilde{\kappa}}{\epsilon^{3}}\varsigma\varrho^{t/2}\lVert\theta-\theta^{\prime}\rVert$ (91)
$\leq 2\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\varsigma\varrho^{t/2}\lVert\theta-\theta^{\prime}\rVert.$ (92)
@@
$\lVert T_{\theta}\tilde{H}_{\theta}(z)-T_{\theta^{\prime}}\tilde{H}_{\theta^{\prime}}(z)\rVert$ $=\left\|\sum_{s=1}^{\infty}(T_{\theta}^{s}H_{\theta}(z)-h(\theta))-\sum_{s=1}^{\infty}(T_{\theta^{\prime}}^{s}H_{\theta^{\prime}}(z)-h(\theta^{\prime}))\right\|$ (93)
$\leq\sum_{s=0}^{\infty}\lVert\tilde{T}_{\theta}^{s}H_{\theta}(z)-\tilde{T}_{\theta^{\prime}}^{s}H_{\theta^{\prime}}(z)\rVert$ (94)
$\leq\left(\frac{\tilde{\kappa}}{\epsilon^{3}}+\frac{\tilde{\kappa}^{2}}{\epsilon^{6}}\right)\frac{2\varsigma}{1-\sqrt{\varrho}}\lVert\theta-\theta^{\prime}\rVert.$ (95)
@@
[…]
@@
Appendix B ESS improvement for the model in Section 5.1
Figure 8 shows how the effective sample size (ESS, Liu,, 1996) of the particle cloud improves while $r_{\lambda}$ is being learned in univariate linear Gaussian model of Section 5.1, to finally reach the performance of the optimal proposal. This is evident for $S_{v}=0.2$; on the other hand, when $S_{v}=1.2$, the particles are well propagated into regions of non-negligible likelihood even with the bootstrap proposal, resulting in the normalized ESS being close to one regardless.
Appendix C Details on the deep generative model of Section 5.3
In this section we provide more details on the model presented in Section 5.3.
C.1. Data generation
@@ C.2. Model architecture
$p(x_{1:T},h_{0:T},y_{1:T})=p_{0}(h_{0})\prod_{t=0}^{T-1}m_{\theta}(x_{t+1}\mid h_{t})g_{\theta}(y_{t+1}\mid h_{t},x_{t+1})p_{\lambda,\theta}(h_{t+1}\mid h_{t},x_{t+1},y_{t+1}),$ (97)
@@
$H_{0}$ $\sim N(0,I),$ (98)
$X_{t+1}\mid H_{t}=h_{t}$ $\sim N(\mu_{\theta}^{x}(h_{t}),\sigma_{\theta}^{x}(h_{t})^{2}),$ (99)
$Y_{t+1}\mid H_{t}=h_{t},X_{t+1}=x_{t+1}$ $\sim\operatorname{Bernoulli}(\mu_{\theta}^{y}(\varphi_{\theta}^{x}(x_{t+1}),h_{t})),$ (100)
$H_{t+1}\mid H_{t}=h_{t},X_{t+1}=x_{t+1},Y_{t+1}=y_{t+1}$ $\sim\delta_{\texttt{GRU}_{\lambda}(h_{t},\varphi_{\theta}^{x}(x_{t+1}),\varphi_{\lambda}^{y}(y_{t+1}))},$ (101)
@@
$X_{t+1}\mid Y_{t+1}=y_{t+1},H_{t}=h_{t}\sim N(\mu_{\lambda}^{p}(\varphi_{\lambda}^{y}(y_{t+1}),h_{t}),\sigma_{\lambda}^{p}(\varphi_{\lambda}^{y}(y_{t+1}),h_{t})^{2}).$ (102)
@@
19:  return $(\xi_{t-38:t+1}^{x,i},\xi_{t-39:t}^{h,i},\omega_{t+1}^{i})_{i=1}^{N},\theta_{t+1},\lambda_{t+1}$.